Aloha Inc. has 6 percent coupon bonds on the market that have 8 years left to maturity. If the YTM on these bonds is 7.5 percent, what is the current bond price? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Annual coupon=1000*6%=60

Hence current price=Annual coupon*Present value of annuity factor(7.5%,8)+$1000*Present value of discounting factor(7.5%,8)

=60*5.857303555+1000*0.560702233

=$912.14(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=$60[1-(1.075)^-8]/0.075

=60*5.857303555

2.Present value of discounting factor=1000/1.075^8

=1000*0.560702233